### v1

# From Flat to Feeling: A Feasibility and Impact Study on Dynamic Facial Emotions in AI-Generated Avatars

Abstract
Dynamic facial emotion is essential for believable AI-generated avatars; however, most systems remain visually inert, limiting their utility in high-stakes simulations such as virtual training for investigative interviews with abused children. We introduce and evaluate a real-time architecture fusing Unreal Engine 5 MetaHuman rendering with NVIDIA Omniverse Audio2Face to translate vocal prosody into high-fidelity facial expressions on photorealistic child avatars. We implemented a distributed two-PC setup that decouples language processing and speech synthesis from GPU-intensive rendering, designed to support low-latency interaction in desktop and VR environments. A between-subjects study ($N=70$) using audio+visual and visual-only conditions assessed perceptual impacts as participants rated emotional clarity, facial realism, and empathy for two avatars expressing joy, sadness, and anger.
Results demonstrate that avatars could express emotions recognizably, with sadness and joy achieving high identification rates. However, anger recognition significantly dropped without audio, highlighting the importance of congruent vocal cues for high-arousal emotions. Interestingly, removing audio boosted perceived facial realism, suggesting that audiovisual desynchrony remains a key design challenge. These findings confirm the technical feasibility of generating emotionally expressive avatars and provide guidance for improving non-verbal communication in sensitive training simulations.
Keywords Human-Computer Interaction; Interactive Avatars; Emotionally Expressive Avatars; Perceived Realism, Unreal Engine MetaHuman; NVIDIA Audio2Face
1 Introduction
AI-generated avatars are rapidly becoming integral components of digital interaction across diverse domains, with applications spanning virtual, augmented, and mixed reality (VR/AR/XR) platforms, virtual telepresence, embodied digital assistants, e-commerce, entertainment, education, and potential therapeutic interventions. The efficacy and adoption of these avatars critically depend on their perceived believability and their capacity to engage users in naturalistic interactions. For an avatar to "feel alive" and effectively bridge the human-computer interaction gap, it must convincingly replicate key facets of human social signaling [1]. Among these, the dynamic expression of facial emotions stands out as particularly vital. Facial expressions serve as a primary, often subconscious, channel for non-verbal communication, conveying a wealth of affective information that shapes interpersonal understanding, relationship dynamics, and the interpretation of intent [2]. Consequently, an avatar’s ability to display appropriate, recognizable, and dynamically evolving facial emotions is paramount for creating engaging, trustworthy, and effective virtual agents. This paper investigates the technical feasibility of creating such avatars and evaluates their perceptual impact, particularly in training scenarios.
Figure 1: Example frames of the two avatars used in the user study: "Emory" (male-presenting) and "Amelia" (female-presenting). The top row displays their neutral facial expressions. The bottom rows show expressions of joy (left) and sadness (right) for each avatar, as used in the study.
While the importance of such affective realism is broadly recognized, its consistent and effective implementation is foundational and challenging in high-stakes domains. One such domain is simulation-based training for professionals who conduct forensic interviews with children, often concerning experiences of abuse or trauma. In this context, previous virtual-child systems have primarily focused on dialogue flow and question typology. However, empirical evaluations have repeatedly highlighted two significant shortcomings [3]: firstly, the avatars themselves are frequently emotionally inert, their faces remaining neutral regardless of the conversational context, thereby offering no micro-expressions that might signal anxiety, relief, or distress. Secondly, practitioners often find the overall visual fidelity of these avatars inadequate, with subtle inconsistencies in proportions, shading, or motion disrupting immersion and undermining trust in the simulation. Collectively, these deficits impede core learning objectives; interviewers are unable to accurately judge whether a question soothes or unsettles the virtual child, cannot adequately fine-tune their empathic responses, and consequently struggle to transfer refined interviewing skills to live cases where non-verbal cues are indispensable. Furthermore, the capacity to render these digital characters in real time is essential for such interactive applications, facilitating immediate responsiveness and a heightened sense of presence.
Addressing these limitations requires a focused research endeavor into developing AI avatar systems capable of rich, dynamic facial emotion expression with real-time responsiveness. This research directly tackles this objective by introducing and assessing a novel, distributed architecture engineered for high-fidelity emotional expressiveness. This system represents a significant advancement by integrating the sophisticated rendering capabilities of Unreal Engine 5 (UE5) MetaHuman software [4] with the intelligent, audio-driven facial animation provided by NVIDIA’s Omniverse Audio2Face [5]. The system’s distributed architecture employs two synchronized computational nodes to effectively balance processing demands—decoupling language understanding and speech generation from GPU-intensive rendering—thereby enabling low-latency interaction. Figure 1 displays representative screenshots of the two photorealistic MetaHuman avatars, "Amelia" and "Emory," utilized in this study.
In summary, our main contributions are as follows:
  • Designing and implementing a modular pipeline for generating emotionally expressive avatars, emphasizing real-time facial animation using NVIDIA Audio2Face and Unreal Engine MetaHuman.
  • Conducting a user study ($N=70$) comparing audio+visual and visual-only conditions, highlighting trade-offs between cross-modal interference and perceived realism.
  • Demonstrating that high-arousal emotions (e.g., anger) rely heavily on audio-visual congruence, while low-arousal emotions (e.g., sadness) achieve robust recognition through visual cues alone.
  • Quantifying how facial structure (angular vs. soft) shapes emotion perception, guiding context-specific avatar design.
  • Identifying key challenges in avatar design and implementation, such as synchronization, the uncanny valley, and expression subtlety, through qualitative feedback analysis.
2 Related Work
The pursuit of emotionally expressive AI avatars lies at the confluence of affective computing, human-computer interaction (HCI), and advancements in multimodal systems, all striving to endow machines with a semblance of human emotional intelligence [6, 7, 8, 9]. While foundational work in affective computing centered on the detection and modeling of emotional cues, the emphasis has progressively shifted towards the dynamic and adaptive expression of these emotions in avatars to foster user engagement and perceived believability. This aligns with broader HCI principles where anthropomorphism in digital agents consistently correlates with enhanced user engagement, perceived credibility, and overall satisfaction [10, 11, 12, 13]. The significance of expressive capabilities is underscored by research in applied domains like virtual therapy and education, where emotionally inert or insufficiently expressive avatars can lead to misunderstandings, diminished rapport, and ultimately, reduced system effectiveness [14, 15, 16, 17].
This need for dynamic emotional expressiveness becomes particularly acute in high-stakes interaction scenarios, such as the virtual training environments for child forensic interviews. In these contexts, an avatar’s ability to display contextually appropriate and recognizable emotions can profoundly influence user perception and training outcomes. However, existing virtual-child systems tailored for such training have predominantly concentrated on dialogue management and question typologies. Empirical evaluations of these systems have consistently identified critical shortcomings: specifically, avatars often lack dynamic facial affect, remaining emotionally static, and their overall visual fidelity is frequently judged as inadequate [3]. These deficiencies impair training efficacy, as trainees cannot accurately gauge a child’s non-verbal reactions to their questioning or refine their empathic responses, which are skills paramount in real-world interviews.
The technological evolution towards more expressive avatars has been considerable. Initial efforts often involved 2D video-based avatars that, while capable of displaying emotions, tended to produce rigid outputs through direct mappings from input signals [18, 19, 20]. The advent of 3D parametric models, such as FLAME [21], offered more granular control over facial geometry and pose. Subsequently, sophisticated speech-driven animation systems like EMOTE [22] and EmoTalk [23] have emerged, aiming to disentangle and map speech content and prosody to facial expressions. However, as these systems push the boundaries of realism, they often encounter challenges in achieving both high-fidelity emotional rendering and the stringent real-time performance demanded by interactive, immersive simulations. Many existing solutions struggle to seamlessly integrate state-of-the-art rendering quality with low-latency, audio-driven emotional animation, particularly for photorealistic child avatars where subtlety and believability are crucial.
Collectively, the literature highlights a persistent gap: the need for AI avatar systems that not only generate rich and dynamic facial emotions but do so with high visual fidelity and in real-time, especially for demanding applications like child interview simulations. While advancements in speech-driven animation and realistic rendering are notable, their integration into a cohesive, responsive system that meets these multifaceted requirements remains a significant hurdle. Our research directly confronts this challenge by proposing and evaluating a novel, distributed architecture. This system uniquely fuses the high-fidelity rendering capabilities of Unreal Engine 5’s MetaHuman with NVIDIA Omniverse Audio2Face for real-time, prosody-driven facial emotion on photorealistic child avatars. By decoupling language processing and speech generation from GPU-intensive rendering via a two-PC setup, our approach is specifically designed to achieve the low-latency interaction essential for effective and immersive training simulations.
Figure 2: Architecture for real-time emotionally expressive child avatars. PC1 processes speech input (Google Cloud STT), generates a text response via a large language model (LLM), and synthesizes speech using Amazon Polly TTS. The resulting audio is analyzed by NVIDIA Audio2Face to extract emotional parameters for facial animation. PC2 renders a customizable Metahuman avatar (Unreal Engine) driven by this emotional data in real time. Synchronization is achieved using the Live Link plugin, and the system supports deployment in both desktop and VR environments.
3 Methodology
3.1 System Design and Implementation
To overcome the limitations of previous systems [3] for emotional expression in AI avatars, this research introduces a novel architecture that integrates Unreal Engine 5 Metahuman [4] for high-fidelity facial rendering with NVIDIA’s Omniverse Audio2Face [5] for generating emotionally expressive facial animations from audio input. Below, we provide a brief overview of these two core platforms:
  • Unreal Engine 5 MetaHuman integrates advanced real-time graphics technologies such as Nanite geometry, the Lumen lighting system, and optimized post-processing to deliver film-quality visuals without compromising interactivity [24, 25]. Using the MetaHuman Creator, researchers can import head scans or digital sculpts and generate fully rigged characters that are customizable in every anatomical detail and exportable at multiple levels of detail (LOD) for deployment across platforms ranging from mobile devices to virtual reality.
    The system includes a preconfigured facial rig built on UE5’s Control Rig, which offers hundreds of easily retargetable controls. This allows expressions to be driven by live capture, audio input, or keyframe animation. To achieve photorealistic skin, which is essential for avoiding the uncanny valley, MetaHuman incorporates subsurface scattering (SSS) [26], a rendering method that simulates the way light penetrates and diffuses through real skin. Additional realism is provided through groom-based hair and physically accurate materials. Together, these features enable a fast end-to-end pipeline for producing responsive and emotionally expressive avatars designed for AI-driven training scenarios involving child interactions.
  • NVIDIA Omniverse Audio2Face is the core component responsible for generating lifelike facial animations in response to audio input. This generative AI tool analyzes speech to automatically produce synchronized lip movements and expressive facial animations in real time, thereby enhancing the emotional realism of virtual characters. It includes an intuitive character retargeting pipeline that enables users to apply animations to custom 3D characters, including seamless integration with Unreal Engine MetaHumans.
    The system provides a Blendshape Generation widget that converts audio into facial expressions using templates such as NV 46 Pose or ARKit 51 Pose. Although the tool primarily supports full-face animation, it can also transfer movements to individual facial features such as the eyes, jaw, and tongue.
    A key advantage of Audio2Face is its real-time performance, which makes it particularly well-suited for interactive applications including virtual reality, digital assistants, and serious games. It also supports traditional animation workflows via live playback or baked outputs. Additional features include multilingual input support, simultaneous animation of multiple characters, and manual adjustment of emotion intensity using facial action unit (FAU) controls. The interface also allows users to fine-tune automatic emotion extraction from audio using settings such as emotion detection range, strength, contrast, smoothing, and the number of concurrently active emotions. These capabilities enable fine-grained, context-aware emotional expression, which is especially important in sensitive domains such as child interview simulations.
The proposed architecture, illustrated in Figure 2, enables real-time interactivity through the coordinated operation of two synchronized computational nodes. PC1 handles all speech and language processing. The user’s spoken input is first transcribed using the Google Cloud Speech-to-Text (STT) engine [27], which leverages advanced models such as Google’s foundation model Chirp to provide high accuracy across diverse accents and acoustic environments, with support for multilingual input. The resulting transcription is processed by OpenAI’s ChatGPT Large Language Model (LLM) [28], which generates a context-appropriate textual response. This response is then synthesized into speech using Amazon Polly’s Text-to-Speech (TTS) system [29], which offers lifelike multilingual voices and allows customization of intonation, pronunciation, and speaking style.
In parallel, NVIDIA Audio2Face analyzes the prosodic and spectral features of the synthesized speech [30]. It extracts frame-level emotional parameters, such as indicators of joy, fear, or sadness, based on vocal characteristics. These parameters are typically represented as blendshape weights or facial action unit (FAU) activations and are transmitted via Wi-Fi to PC2, the rendering node.
PC2 is responsible for avatar rendering and real-time facial animation. It runs Unreal Engine 5’s Metahuman framework, which produces high-fidelity child avatars with anatomically accurate facial rigs. The emotional parameters received from PC1 are used to drive the avatar’s expressions in real time by mapping the blendshape weights or FAU activations to the MetaHuman Control Rig via the Live Link plugin [31]. This ensures low-latency synchronization between facial animation and the synthesized voice.
This distributed system design separates the computationally intensive processes of natural language understanding and speech generation from the demanding task of real-time, high-resolution avatar rendering. While this architecture is designed for responsiveness, potential sources of audiovisual desynchrony can still arise from factors such as network latency or processing delays within either computational node. Nevertheless, this separation enables the system to maintain both overall responsiveness and visual realism, supporting immersive and emotionally expressive interactions in applications such as child interview training.
3.2 User Study
To examine the feasibility of incorporating emotional expressiveness into AI-driven avatars through dynamic facial emotion expression and to assess its perceptual impact, we conducted a user study focused on three key dimensions: emotional clarity, facial realism, and empathic connection. These dimensions were chosen as they are key indicators of avatar believability and effectiveness in interactive simulations. Two photorealistic MetaHuman avatars, named Emory and Amelia, were used to assess perceived emotional expressiveness under different conditions. Representative screenshots of the avatars are shown in Figure 1.
The study followed a between-subjects design with two experimental conditions: one condition included both voice and facial expression (audio+visual condition), while the other featured facial expression only (visual-only condition). This design was informed by prior findings suggesting that audio and visual modalities contribute differently to emotion perception, specifically that audio cues are more salient for detecting changes in emotional arousal, while visual cues are more informative for identifying emotional valence [32].
A total of 70 participants were recruited via the Prolific platform and randomly assigned to one of two experimental conditions ($N=35$ per condition). The sample exhibited diversity in gender, age, academic/professional background, and familiarity with virtual environments.
The participant pool ($N=70$) was evenly distributed by gender, comprising 35 women (50.0%) and 35 men (50.0%). The predominant age group was 25–34 years (40 participants, 57.1%), followed by 35–44 years (16 participants, 22.9%). Smaller segments included participants aged 18–24 years (10 participants, 14.3%) and 45–54 years (4 participants, 5.7%). The academic and professional backgrounds of the participants were highly varied, with numerous unique responses.
Regarding familiarity with virtual environments and avatars, a significant portion of participants reported interacting "Occasionally" (32 participants, 45.7%). Other interaction frequencies included "Rarely" (15 participants, 21.4%), "Frequently" (12 participants, 17.1%), and "Very frequently" (9 participants, 12.9%). A small fraction (2 participants, 2.9%) reported "Never" interacting with avatars. This demographic profile indicates a sample primarily composed of young to middle-aged adults from diverse professional domains, possessing a general, though not necessarily expert-level, familiarity with avatar interactions.
Participants viewed a series of six high-definition video clips, the presentation order of which was counterbalanced to avoid order effects. In these clips, the avatars expressed three emotional states: sadness, anger, and joy (see Figure 3). These specific emotions were selected as they represent a range of valence and arousal levels and are relevant emotions that children might express in sensitive interview contexts. In the audio+visual condition, speech was synthesized using OpenAI’s TTS system to ensure consistent vocal tone and emotionally rich delivery. Since child or adolescent voices were not available, we used a young adult female voice for both avatars. This choice was supported by findings from our prior studies [33], which showed that female voices can be effectively used to represent child characters in training scenarios, regardless of gender.
Sadness "I thought today would be different… but it wasn’t. No one talked to me at school. I just sat alone… again. I don’t know why they don’t like me. Maybe I’m just not good enough."
Anger "It’s not fair! I told them it wasn’t my fault, but no one listened. They just blamed me like always! I’m so tired of this. I try to do the right thing, but it never matters!"
Joy "Today was amazing! My teacher said I did a great job on my homework, and my friends invited me to play during lunch. Everything just felt really good … like things are finally going my way!"
Figure 3: Scripted utterances used for emotional avatar expressions in the user study.
The six video clips were generated using the described system architecture. For the audio+visual condition, the scripts were synthesized into audio, and this audio was processed by NVIDIA Audio2Face to generate facial animation data. This animation data was then applied to the MetaHuman avatars in Unreal Engine 5 and rendered into video clips. For the visual-only condition, the same facial animation data was applied to the avatars and rendered into video clips without the synthesized audio track.
To manage participant expectations and establish a perceptual anchor, we began with a deliberately low-fidelity stimulus, following evidence that anchoring effects require the presence of both semantic and numerical cues [34]. The inclusion of this initial low-fidelity stimulus was intended to provide a baseline or reference point for participants, potentially making their subsequent ratings of the higher-fidelity main stimuli more calibrated and less susceptible to being skewed by initial unfamiliarity or overly high expectations for avatar realism. By first exposing them to a clearly less sophisticated avatar, we aimed to help participants better contextualize and evaluate the nuances of the emotionally expressive avatars that formed the core of the study.
After viewing the main stimuli, participants completed a questionnaire comprising three main sections (see Table 1). Most items were rated on a five-point Likert scale and assessed aspects such as emotional clarity (how well the intended emotion was recognized), realism of facial expressions, and perceived emotional connection or empathy felt toward the avatars. The final section included an open-ended question inviting participants to reflect on what stood out most about the avatars.
Table 1: Structure and Content of the User Study Questionnaire
Section Question Type Content
Emotion Recognition Likert scale Extent to which the avatar expressed: Anger, Sadness, Joy, Fear, Disgust
Facial Realism Likert scale Realism of facial expressions in each video
Realism Contributors Likert scale Contribution of each factor to perceived realism:
- Facial expressions
- Visual appearance
- Voice tone*
- Dialogue content*
Empathy/Connection Likert scale Emotional connection or empathy felt toward the avatars overall
Open-ended Reflection Free text One sentence describing what stood out most about the avatars
Demographics Multiple choice Gender, age group, academic/professional background, frequency of avatar interaction
* Voice tone and dialogue content were excluded from the visual-only condition.
4 Results and Analysis
This section presents the findings of the user study, analyzing how participants perceived emotional expressions across different conditions and avatars, based on both quantitative ratings and qualitative feedback.
To evaluate perceptual differences across experimental conditions and avatars, statistical comparisons were performed using independent-samples $t$-tests or Welch’s $t$-tests. The specific test was chosen based on the results of Levene’s test for equality of variances. This approach ensures statistically robust comparisons while appropriately accounting for potential heteroscedasticity, as recommended by Field et al. [35]. The outcomes of these comparisons are detailed in Table 2 (for audio+visual vs. visual-only conditions), Table 3 (for general questionnaire items), and Table 4 (for comparisons between the two avatars).
Descriptive statistics, specifically means ($M$) and standard deviations ($SD$), were calculated for each variable under investigation. These descriptive data are graphically visualized in Figure 4. For enhanced clarity and direct comparison, all aforementioned tables (Table 2, Table 3, and Table 4) also present these mean values alongside the corresponding inferential test statistics, such as $t$-statistics and $p$-values.
4.1 Comparison of Audio and Visual Versus Visual-Only Conditions
Figure 4: Mean ratings ($\pm$±SD) of emotional expressiveness across five emotions (Angry, Sad, Joy, Fear, Disgust) for the avatars Emory and Amelia, under both audio+visual and visual-only conditions. Each video clip is labeled on the left, with color-coded bars distinguishing the experimental conditions. Green-highlighted boxes indicate intended emotions whose recognition was especially important for the study.
[…]
Variable t-statistic p-value
Mean
Audio+Visual
Mean
Visual-Only
Facial exp. 2.843 0.006 2.800 3.429
Angry -2.405 0.019 2.571 1.829
Sad -0.923 0.360 3.629 3.343
Joy 1.579 0.119 1.171 1.429
Fear 1.318 0.192 1.857 2.229
Emory_Sad Disgust 0.000 1.000 1.971 1.971
Facial exp. 2.239 0.028 2.914 3.514
Angry -10.729 0.000 3.829 1.429
Sad -4.122 0.000 3.257 2.000
Joy 5.559 0.000 1.086 2.143
Fear -0.762 0.448 1.914 1.714
Emory_Angry Disgust -4.306 0.000 2.743 1.486
Facial exp. 1.222 0.226 3.171 3.514
Angry 2.766 0.008 1.057 1.400
Sad 3.000 0.005 1.114 1.743
Joy -2.887 0.005 3.771 2.914
Fear 1.768 0.082 1.229 1.571
Emory_Joy Disgust 2.743 0.009 1.171 1.743
Facial exp. 2.836 0.006 3.343 4.000
Angry -0.803 0.425 2.571 2.314
Sad -0.125 0.901 4.114 4.086
Joy 1.226 0.224 1.086 1.200
Fear 0.751 0.455 2.086 2.314
Amelia_Sad Disgust -2.145 0.036 2.429 1.800
Facial exp. 0.351 0.726 3.543 3.629
Angry -2.988 0.004 3.114 2.171
Sad -4.172 0.000 3.714 2.371
Joy 2.928 0.005 1.143 1.829
Fear -0.543 0.589 2.000 1.857
Amelia_Angry Disgust -0.560 0.577 2.114 1.943
Facial exp. 1.680 0.098 3.486 3.914
Angry 1.358 0.179 1.000 1.086
Sad 0.848 0.400 1.057 1.143
Joy 0.838 0.405 4.200 4.371
Fear 0.314 0.754 1.114 1.143
Amelia_Joy Disgust 1.577 0.119 1.057 1.229
The analysis of perceptual differences between the audio+visual and visual-only conditions is structured as follows:
  • Perceived Facial Expressions Realism: As shown in Table 2, a consistent pattern emerges in participants’ ratings of facial expression realism across conditions. Participants rated the visual realism of facial expressions significantly higher in the visual-only condition across all six stimulus clips. Mean scores for "Facial expressions Realism" uniformly increased when auditory input was removed. Statistically significant differences were observed in three of the six clips: Emory-Sad ($p=0.006$), Emory-Angry ($p=0.028$), and Amelia-Sad ($p=0.006$). This phenomenon likely arises from audiovisual interference; the presence of synthetic speech may divide attentional resources between auditory and visual channels, thereby amplifying minor artifacts such as lip desynchronization. In the visual-only condition, the unimodal evaluation eliminates such cross-modal incongruities, allowing participants to assess facial animations in isolation, which consequently enhances perceived realism.
  • Recognition of Sadness and Joy: Findings presented in Table 2 and Figure 4 indicate the emotions of sadness and joy were consistently recognized across both experimental conditions, with both avatars achieving near-ceiling recognition scores for these emotions. For Amelia, sadness recognition remained stable between the audio+visual ($M=4.11$) and visual-only ($M=4.09$) conditions. Her joy recognition scores showed a slight, non-significant increase in the visual-only condition ($M=4.37$) compared to the audio+visual condition ($M=4.20,p=0.41$). In contrast, Emory’s sadness recognition showed minimal variation between the audio+visual ($M=3.63$) and visual-only ($M=3.34$) conditions. However, his joy recognition declined sharply and significantly in the visual-only condition, dropping from $M=3.77$ to $M=2.91$ ($p=0.005$).
    Sadness recognition remained stable across conditions, consistent with its reliance on visual cues like brow depression and lip-corner droop (FACS) [36]. Joy recognition, however, diverged by avatar: Amelia’s scores were robust in both conditions, likely due to clear Duchenne markers (e.g., cheek elevation), while Emory’s joy recognition dropped 22% in visual-only condition. This suggests participants relied on auditory prosody (e.g., rising pitch) to distinguish Emory’s joy from frustration; in the visual-only condition, his brow tension was misinterpreted as negative affect, underscoring the interplay between modality and avatar morphology. These findings confirm that sadness relies on visual fidelity, while joy benefits from audiovisual congruence, particularly through Duchenne smiles. Auditory cues minimally affected sadness but were critical for disambiguating Emory’s joy, aligning with prior work on facial action units (AUs) in low-arousal emotions [2]. The variance in joy recognition highlights modality’s context-dependent effects. Amelia’s visual clarity reduced cross-modal interference, improving joy perception without audio. Conversely, Emory’s angular morphology required vocal reinforcement to resolve ambiguity, exposing risks in incongruent multimodal design. This asymmetry challenges universal design rules, advocating instead for emotion-specific optimization: prioritize audiovisual synchronization for high-arousal emotions (e.g., joy) and visual precision for low-arousal ones (e.g., sadness). Such tailored approaches could mitigate mismatches between morphology and acoustic cues, advancing emotionally intelligent avatar design.
  • Recognition of Anger: Table 2 and Figure 4 show that anger recognition most strongly relied on audio cues for both avatars. Emory’s anger scores dropped from $M=3.83$ (audio+visual) to $M=1.43$ (visual-only), and Amelia’s from $M=3.11$ to $M=2.17$. These significant declines (Table 2; Emory: $p<0.001$, Amelia: $p=0.004$) accompanied frequent misclassifications of silent anger displays as mild sadness, with brow furrows seen as worry, not aggression. This aligns with appraisal theory and acoustic studies, indicating high-arousal emotions like anger depend on prosodic cues (e.g., vocal intensity, spectral noise). Without these, facial scowls alone were insufficient for disambiguation, stressing the need for multimodal congruence. Notably, even with audio, Amelia’s anger ($M_{anger}=3.11$) was perceptually overshadowed by sadness ($M_{sadness}=3.71$), unlike Emory’s anger-dominant profile ($M_{anger}=3.83$ vs. $M_{sadness}=3.26$), suggesting avatar morphology and voice design also modulate emotional salience.
    This perceptual asymmetry for anger likely stems from morpho-acoustic factors. Amelia’s softer features (e.g., jawline, subdued lip compression) may have reduced perceived aggression. Furthermore, her synthesized young female vocal profile lacked typical low-frequency anger cues, potentially biasing interpretations towards sorrow. These results reiterate the need for congruent multimodal signals for high-arousal emotions. Future work should investigate augmenting Amelia’s upper-face tension (e.g., sharper V-brows, AU4 activation) and integrating harsher prosody to enhance anger clarity without sacrificing fidelity.
  • Perceived realism and Empathy levels: The analysis presented in Table 3 explores two key dimensions: firstly, the influence of multimodal (audio and visual) cues on participants’ perception of realism, and secondly, their self-reported levels of empathy.
    In the audio+visual condition, auditory elements emerged as significant contributors to perceived realism. Specifically, voice tone received a mean rating of $4.0$, and dialogue content was rated at $M=3.86$. Conversely, aspects related to visual fidelity—namely facial expression realism and overall visual appearance—were rated higher in the visual-only condition ($M=3.86$ and $M=3.80$, respectively) compared to the audio+visual condition ($M=3.29$ and $M=3.40$, respectively). The difference in perceived facial expression realism between the two conditions was statistically significant ($t=2.21,p=0.03$). This suggests that the exclusion of auditory input may have reduced cross-modal interference, thereby enabling participants to focus more intently on, and assign higher ratings to, the visual fidelity of the avatars.
    When empathy levels were analyzed independently, no statistically significant difference was found between the visual-only condition ($M=3.06$) and the audio+visual condition ($M=2.83$; $t(68)=0.75,p=0.46$). While the means suggest a slight numerical increase in reported empathy in the visual-only condition, this difference did not reach statistical significance. One might speculate that, had a stronger effect been present, it could align with prior research suggesting that a unimodal visual focus might enhance attention to subtle facial expressions, potentially deepening emotional engagement. However, several factors in the current study could have influenced these results. For instance, in the audio+visual condition, the use of a young adult female voice for child avatars (a necessary choice due to the unavailability of suitable child or adolescent synthesized voices) may have created a perceptual disconnect for some participants. Furthermore, any instances where the synthesized voice did not fully align with the intended emotion, or where subtle audiovisual mismatches were perceived, could have potentially detracted from the emotional impact in the audio+visual condition. These factors might have counteracted any benefits of congruent multimodal cues for empathy in this specific experimental setup.
[…]
Variable t-statistic p-value
Mean
Audio+Visual
Mean
Visual-Only
Facial expressions 2.21 0.03 3.29 3.86
Visual appearance 1.39 0.17 3.40 3.80
Voice tone N/A N/A 4.0 N/A
Dialogue content N/A N/A 3.86 N/A
Empathy level 0.75 0.46 2.83 3.06
Ultimately, these results highlight a critical trade-off: while auditory signals are indispensable for effectively conveying anger and resolving ambiguities in high-arousal states, their inclusion risks exposing subtle audiovisual mismatches that can undermine perceived realism. Achieving seamless synchronization between vocal prosody and facial kinematics thus emerges as a prerequisite for maximizing both emotional clarity and immersion in virtual interactions.
4.2 Comparison of Emory Versus Amelia
[…]
Audio+Visual Visual-Only
Variable t-statistic p-value Mean_Amelia Mean_Emory t-statistic p-value Mean_Amelia Mean_Emory
Facial exp. 3.496 0.001 3.343 2.800 3.824 0.001 4.000 3.429
Angry 0.000 1.000 2.571 2.571 1.759 0.088 2.314 1.829
Sad 2.451 0.020 4.114 3.629 2.481 0.018 4.086 3.343
Joy -1.139 0.263 1.086 1.171 -1.604 0.118 1.200 1.429
Fear 1.186 0.244 2.086 1.857 0.502 0.619 2.314 2.229
Sad Disgust 2.604 0.014 2.428 1.971 -1.099 0.280 1.800 1.971
Facial exp. 3.421 0.002 3.543 2.914 0.560 0.579 3.629 3.514
Angry -3.446 0.002 3.114 3.829 2.926 0.006 2.171 1.429
Sad 2.758 0.009 3.714 3.257 1.625 0.113 2.371 2.000
Joy 0.572 0.571 1.143 1.086 -1.429 0.162 1.829 2.143
Fear 0.517 0.609 2.000 1.914 0.709 0.483 1.857 1.714
Angry Disgust -2.666 0.012 2.114 2.743 3.053 0.004 1.942 1.486
Facial exp. 1.995 0.054 3.486 3.171 2.119 0.041 3.914 3.514
Angry -1.435 0.160 1.000 1.057 -2.946 0.006 1.086 1.400
Sad -1.000 0.324 1.057 1.114 -3.340 0.002 1.143 1.743
Joy 2.445 0.020 4.200 3.771 6.065 0.000 4.371 2.914
Fear -1.160 0.254 1.114 1.229 -2.321 0.026 1.143 1.571
Joy Disgust -1.675 0.103 1.057 1.171 -2.927 0.006 1.229 1.743
A comparative analysis of the emotional expressiveness specific to each avatar revealed distinct perceptual differences between Amelia, the female-presenting avatar, and Emory, the male-presenting avatar (Table 4). Participants consistently assigned higher ratings to the realism of Amelia’s facial expressions across both audio+visual and visual-only conditions. Statistically significant differences (highlighted in gray in Table 4) emerged in four of the six emotion-condition pairings: sadness in both the audio+visual ($p=0.001$) and visual-only ($p=0.001$) conditions, anger in the audio+visual condition ($p=0.002$), and joy in the visual-only condition ($p=0.041$). These findings indicate a consistent perceptual preference for Amelia’s facial animations across various emotional contexts and presentation modalities, a preference likely influenced by her morphological design.
One potential explanation for Amelia’s superior expressiveness lies in her softer facial geometry and the subtlety of her micro-expressions, which appeared to align more closely with viewers’ expectations of emotional authenticity. For instance, her rounded jawline and gentle brow dynamics effectively amplified the perceived realism of sadness (e.g., downturned lips) and joy (e.g., symmetrical smiles). This observation resonates with previous research on anthropomorphic design, which suggests a correlation between such features and perceived warmth [37].
However, this pattern diverged in the context of anger recognition, where Emory demonstrated superior performance ($p=0.002$). Amelia’s more subdued brow contraction and reduced lip tension appeared to attenuate the visual cues typically associated with aggression. Consequently, in the visual-only condition, 43% of participants misclassified Amelia’s expression of anger as sadness. This finding is consistent with appraisal theory, which underscores the importance of congruent visual and auditory signals for disambiguating high-arousal emotions [38, 39]. In contrast, Emory’s angular facial features and more pronounced brow kinematics served to amplify the perception of aggression. This highlights a potential design trade-off: while a subtler morphology may enhance the realism of low-arousal emotions, it risks ambiguity when portraying high-arousal states. This trade-off is also reflected in findings from multimodal deep learning frameworks, which indicate that the joint modeling of auditory and visual cues improves emotion recognition accuracy, particularly for high-arousal emotions like anger [40]. Furthermore, these perceptual divergences align with cross-cultural research on the interpretation of facial expressions, which demonstrates that morphological features and subtle dynamic expressions can significantly influence emotion attribution. Such research shows that specific facial movements may be interpreted differently depending on both cultural background and structural facial cues [41].
4.3 Qualitative Feedback
Participants provided reflective qualitative feedback on what stood out the most about the avatars. These reflections revealed distinct and important observations regarding emotional expressiveness, realism, and immersion, providing directions for enhancing avatar technologies in interactive simulations.
4.3.1 Audio and Visual Conditions
In the audio+visual condition, participants’ qualitative feedback highlighted both the strengths and limitations of the avatars’ emotional expressiveness. On the positive side, many participants recognized and appreciated the avatars’ capacity to convey emotions dynamically, noting that "They were all expressing emotions" and that they "expressed their feelings through facials". The quality of the synthesized voice was frequently mentioned as a significant contributor to engagement and realism, with comments such as "The tone of voice was the most convincing" and "The dialogue content… did not sound like an avatar", suggesting successful prosody design and auditory fidelity. Some participants also felt a sense of connection, stating, "You can feel them though facial expressions". The visual appearance was also praised, with feedback indicating that "Their visual appearances were very good and looked human like" and "The faces were well done".
However, a recurring theme in the feedback was the presence of multimodal incongruities. Participants frequently observed a mismatch between the auditory and visual channels, as captured by comments like "The facial expression does not match what the avatars are saying". Issues with lip synchronization were also noted as detrimental to realism, with feedback stating, "When they speak, their lips do not correspond to the dialogue" and "The mouth moves but it doesn’t seam connected to the rest of the face". These observations align with the quantitative findings regarding reduced perceived facial realism when audio was present (Table 2), suggesting that the inclusion of voice amplified scrutiny of temporal and spatial alignment.
Participants also reported experiencing uncanny valley effects, describing the avatars with phrases such as "weird mix of looking real but still obviously fake" and comparing their movement to "NPC characters in video games" or "a rigid model in video games". Some felt the facial expressions were not consistently on point or that the faces "didnt move all that much". Negative emotional responses were also reported by some participants, with comments like "They all looked very scary" and "I feel uncomfortable and uneasy after this".
Collectively, the feedback from the audio+visual condition reveals a tension between the successful implementation of some aspects, such as voice quality and general emotional expression, and persistent challenges related to audiovisual synchronization and the subtlety and naturalness of facial movements. These points underscore the complexity of creating truly believable and comfortable AI-generated avatars.
4.3.2 Visual-Only Conditions
In the visual-only condition, where avatars were presented without synthesized speech, participant feedback predominantly focused on visual realism and the effectiveness of facial cues in conveying emotion. Positive comments highlighted the avatars’ ability to express emotions through isolated facial dynamics, with participants noting that "Facial expressions stood out the most from the avatars" and that the avatars showed emotion "like real human". Some feedback specifically praised the clarity of certain expressions, such as joy and sadness, with one participant observing a "genuine smile, often characterized by crinkles around the eyes" for happiness and another emphasizing the effectiveness of "The emotional that was in the eyes of avatars… specially the emotion of sadness and disgust," adding "good job!". The visual design was also acknowledged positively, with comments like "They have human-like features" and "Amelia and Emory looked the most realistic to me". Some participants reported feeling "Very connected with the avatar" and that "You can feel them though facial expressions" even in the absence of audio.
However, critiques were also present, centering on perceived exaggeration or inconsistency in the expressions. Participants reported that "facial expressions were quite exaggerated and did not look completely natural" and that "They almost all seem to have the same facial expression if I did not look closely". Technical limitations in animation fidelity were also noted, with comments such as "I think they lack polygons to make them more expresive" and comparisons to "a rigid model in video games". The absence of audio seemed to shift attention to subtle visual flaws, including repetitive mouth movements ("the movement of their mouth") and limited emotional range or gradation ("I would see a slight smile here and there but the avatar still kept a straight face"). Some feedback also simply stated, "nothing" stood out, or described the experience as "interesting but not very realistic".
Collectively, the feedback from the visual-only condition suggests that while removing audio can enhance the focus on and perceived realism of facial expressions by eliminating cross-modal interference, it also exposes limitations in the subtlety, naturalness, and fidelity of the facial animations themselves. These points underscore the need for further refinement in visual-only emotional expression to achieve greater believability.
5 Discussion
This research successfully demonstrates the technical viability of a distributed architecture that integrates Unreal Engine 5’s MetaHuman and NVIDIA Omniverse Audio2Face for generating real-time, high-fidelity emotional expressions in AI-driven child avatars. Our empirical evaluation yielded several critical insights. Firstly, the system can synthesize recognizable emotional states, with low-arousal emotions like sadness and joy achieving high identification rates irrespective of whether audio was present. Secondly, the accurate recognition of high-arousal emotions, notably anger, is significantly dependent on congruent vocal cues, with performance dropping substantially in visual-only conditions. Thirdly, avatar morphology plays a crucial role in emotion perception; the softer-featured "Amelia" avatar was generally perceived as more lifelike, whereas the more angular "Emory" avatar more effectively conveyed anger, highlighting a potential trade-off between perceived warmth and the unambiguous display of high-arousal states.
A persistent and significant challenge underscored by this study is audiovisual desynchrony. The marked increase in perceived facial realism when audio was removed suggests that even subtle temporal or qualitative mismatches between auditory and visual channels can considerably disrupt the user’s sense of realism. This phenomenon, also noted in participant feedback regarding lip-sync issues and expressions not matching dialogue, remains a key design bottleneck. Achieving seamless synchronization is not merely a technical refinement but a prerequisite for maximizing both emotional clarity and immersion. This difficulty likely stems from the inherent complexities of perfectly aligning machine-generated prosody and facial kinematics with human perceptual expectations, which are highly attuned to subtle inconsistencies.
Furthermore, the choice of generation parameters within the animation tools themselves warrants careful consideration. In this study, an "exaggerated mode" was utilized for the NVIDIA Audio2Face settings. While intended to ensure distinct emotional expression, this may have contributed to perceptions of unnatural or overly pronounced facial movements, as noted by some participants in the visual-only condition. Exploring "automatic" or more nuanced settings could yield more subtle and naturalistic expressions, potentially mitigating these issues and enhancing overall realism. Future work should systematically investigate the impact of these diverse settings. The inherent capabilities of technologies like MetaHuman, particularly features such as subsurface scattering for realistic skin rendering, offer significant potential in overcoming the uncanny valley. However, realizing this potential fully depends on harmonizing these advanced visuals with appropriately matched age-specific voice, precise lip synchronization, and nuanced emotional expressions driven by tools like Audio2Face.
5.1 Limitations
While this study provides key findings, certain limitations should be acknowledged:
  • Avatar Specificity: The findings are based on two specific MetaHuman avatars. Perceptions might differ with avatars of varying design, age representation, or art style.
  • Voice Realism and Appropriateness: The use of a young adult female voice for child avatars, though based on prior work, may have influenced perceptions of authenticity and emotional congruence. The availability and integration of high-quality, age-appropriate synthesized child voices remain a challenge.
  • Emotion Scope: The study focused on three basic emotions (joy, sadness, anger). The system’s efficacy in portraying a wider range or more complex emotional states (e.g., fear, surprise, confusion) was not assessed.
  • Participant Sample: Participants were recruited via an online platform (Prolific). While providing diversity, this sample may not fully represent the target end-users of specialized training systems, such as Child Protection Services (CPS).
  • Controlled Environment: The study was conducted using pre-rendered video clips. Interaction in a fully real-time, dynamic environment might elicit different perceptual responses.
5.2 Future Work
Several avenues for future research emerge from this study. The observation, echoed in user comments, that perceptions and responses may evolve with continued exposure highlights the need for longitudinal studies. Such research would offer valuable insights into adaptation, learning effects, and the long-term engagement potential of emotionally expressive avatars. Developing globally applicable systems necessitates exploring cultural variations in the expression and perception of emotions. Avatar design and emotional synthesis should account for these differences to ensure broader relatability and avoid misinterpretation. Finally, a critical next step is to conduct user studies with domain experts, such as professionals in CPS or law enforcement. Evaluating the effectiveness and usability of these emotionally expressive avatars within realistic child forensic interview training simulations will provide invaluable feedback. This will help identify areas for refinement to meet the specific, nuanced requirements of these high-stakes environments, ultimately assessing the practical impact of the system in its intended application context.
6 Conclusion
This research successfully establishes the feasibility of a novel, real-time architecture integrating Unreal Engine 5 MetaHuman and NVIDIA Omniverse Audio2Face for generating dynamic and nuanced facial emotions in AI-generated avatars. The core achievement lies in demonstrating this system’s capacity to produce recognizable emotional expressions, particularly for low-arousal states like joy and sadness, thereby marking a significant step towards creating more believable and effective virtual agents. These advancements are especially pertinent for high-stakes applications, such as virtual training simulations for child forensic interviews, where emotional realism can profoundly impact training efficacy.
Our findings underscore the differential perceptual impact of emotions based on arousal levels and the crucial role of multimodal congruence, revealing that while visual cues suffice for some emotions, accurately conveying high-arousal states like anger critically depends on synchronized audiovisual input. The study further pinpointed audiovisual synchronization and avatar morphology as pivotal factors influencing perceived realism and emotional clarity, offering key technical and design considerations for future development in the field. While challenges such as navigating the uncanny valley and refining the subtlety of expressions persist, this work clearly demonstrates the considerable potential of combining high-fidelity rendering with sophisticated audio-driven animation. Beyond specialized training, the principles and insights derived contribute to the broader endeavor of creating more naturalistic, engaging, and perceptually rich artificial characters across diverse interactive environments. This study provides a foundational contribution to the growing body of research on emotionally expressive avatars, paving the way for future innovations in human-AI interaction.